Two events  and  associated with a sample space of an experiment are such that       ,  and  . Determine whether   and  are independent.


1
Expert's answer
2021-07-20T04:20:23-0400
"P(E_1\cup E_2)=P(E_1)+P(E_2)-P(E_1\cap E_2)"

"P(E_1\cap E_2)=P(E_1)+P(E_2)-P(E_1\cup E_2)"

"=0.27+0.40-0.58=0.09"

"P(E_1)P(E_2)=0.27(0.40)=0.108"

"\not=0.09=P(E_1\cap E_2)"



Since "P(E_1\cap E_2)\not=P(E_1)P(E_2)," then "E_1" and "E_2" are not independent.